Find [H3O+] if the pH = 1.4.


Solve the problem. Round your answer to the nearest tenth, when appropriate. Use the formula pH = -log[H3O+], as needed. 

Find [H3O+] if the pH = 1.4. 

  • 4.0 x 10-1 

  • 2.5 x 10-1 

  • 4.0 x 10-2 

  • 2.5 x 10-2
Answer #1

We know that

pH = - log[H3O+]

Given, pH = 1.4

=> - log[H3O+] = 1.4

=> log[H3O+] = - 1.4

=> [H3O+] = 10-1.4

=> [H3O+] = 0.04

=> [H3O+] = 4.0 * 10-2

Therefore,

[H3O+] = 4.0 * 10-2